An urn contains four red balls, six black balls, and five green balls. If two balls are selected at random without replacement from the urn, what is the probability that a red ball and a black ball will be selected? (Round your answer to three decimal places.)
Answer #1

Total balls = 4 + 6 + 5 = 15

P(red ball and black ball) = 4C1 * 6C1 / 15C2 = 4 * 6 / 105 = 0.229